ganeshie8 (ganeshie8):

Last question on elasticity A 103 kg uniform log hangs by two steel wires, A and B, both of radius 1.20 mm. Initially, wire A was 2.50 m long and 2.00 mm shorter than wire B. The log is now horizontal. What are the magnitudes of the forces on it from (a) wire A and (b) wire B? (c) What is the ratio dA/dB?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

I know that I need to use young's modulus of steel and the stress-strain relation : \[\dfrac{F}{A} = Y\dfrac{\Delta L}{L}\]

ganeshie8 (ganeshie8):

Before that, balancing vertical forces I get : \[103g = T_A+T_B \tag{1}\]

ganeshie8 (ganeshie8):

I am not sure yet how to make use of that stress-strain relation here...

Parth (parthkohli):

Hmm, OK. I haven't studied this so I'll be counting on you for the most part. If shorter wire stretches by \(\Delta L\) then the longer one must stretch by \(\Delta L -2~\rm mm \) so that the log is horizontal in the end.

ganeshie8 (ganeshie8):

oh, right : \[\Delta L_B = \Delta L_A -0.002\tag{2}\]

ganeshie8 (ganeshie8):

two extra unknowns hmm ..

ganeshie8 (ganeshie8):

From stress strain relation : \[\dfrac{T_A}{\pi(1.2/1000)^2} = 200*10^9*\dfrac{\Delta L_A}{2.5-0.002}\tag{3}\] \[\dfrac{T_B}{\pi(1.2/1000)^2} = 200*10^9*\dfrac{\Delta L_B}{2.5}\tag{4}\]

ganeshie8 (ganeshie8):

four equations and four unknowns : \(T_A, T_B, \Delta L_A, \Delta L_B\) i hope they are independent equations... let me feed them to wolfram :)

ganeshie8 (ganeshie8):

any idea on how to work part c ?

Parth (parthkohli):

Net torque = 0?

ganeshie8 (ganeshie8):

\(-d_AT_A +d_BT_B=0 \implies \dfrac{d_A}{d_B} = \dfrac{T_B}{T_A}\) perfect! thanks

OpenStudy (fairegaming):

Omg so smart, I will never be that good in math Q.q mind blown.